Sunland, Inc., has revenue of $471,000, costs of $329,700, and a tax rate of 31 percent. If the firm pays out 50 percent of its earnings as dividends every year, how much earnings are retained and what is the firm’s retention ratio? (Round answers to 0 decimal places e.g. 1,575.)

Retained earnings $

Retention Rate %
Answer #1

Profit =(Sales-Costs)*(1-Tax Rate) =(471000-329700)*(1-31%) =97497
Earnings Retained =Profit*(1-Pay Out Ratio)=97497*(1-50%) =48748.50
Retention Ratio =1-Pay out Ratio =1-50% =50%